OpenStudy (anonymous):

integration cos^2 theta * sin theta (d theta)

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} \cos^2 \theta \sin \theta d \theta\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hmm

OpenStudy (amoodarya):

hint : cos(th)=u

OpenStudy (anonymous):

i got it! thank you!

OpenStudy (amoodarya):

ok your welcome